# Leibniz rule for wedge product in discrete exterior calculus on general polygonal meshes

### v1

@@ Abstract
Discrete exterior calculus offers a coordinate–free discretization of exterior calculus especially suited for computations on curved spaces. In this work, we present a wedge product on 2–dimensional pseudomanifolds, whose faces are any polygons. We prove that this polygonal wedge product is compatible with the discrete exterior derivative in the sense that it satisfies the Leibniz product rule. We thus extend previously studied discretizations of wedge products from simplicial or quadrilateral meshes to general polygonal surface meshes. We also prove that our discrete wedge product corresponds to a cup product of cochains on 2–pseudomanifolds.
@@
Discrete wedge product, Discrete exterior calculus, Polygonal wedge product, Polygonal cup product, Discrete differential forms, Polygonal exterior calculus
pacs:
MSC Classification]65N22, 57Z20, 57-08, 53Z30
@@ 1 Introduction
Discrete versions of exterior calculus on surface meshes have gained increased attention in the last few decades, see [1, 2, 3, 4] and references therein. Discete exterior calculus (DEC) and finite element exterior calculus (FEEC) have found applications in computational fluid dynamics [5, 6], computational electromagnetism [7], and computer graphics [8, 4]. See also references in [9] for more applications.
In this paper, we prove that the wedge product on general polygonal meshes, first introduces in [10] and later used in applications in [4], satisfies the Leibniz product rule with discrete exterior derivative (coboundary operator). Furthermore, we show that it corresponds to a cup product of cochains on orientable 2–dimensional pseudomanifolds. We thus partially extend the findings of [9] from simplicial to general polygonal setting.
On smooth manifolds, the wedge product allows for building higher degree forms from lower degree ones. Similarly a cup product is a product of two cochains of arbitrary degree $p$ and $q$ that returns a cochain of degree $p+q$ located on $(p+q)$–dimensional cells. The cup product was introduced by J. W. Alexander, E. Čech, and H. Whitney [11] in 1930’s and it became a well–studied notion in algebraic topology, mainly in the simplicial setting [12, 13]. Later, the cup product was extended also to $n$–cubes [14]. In [15] various cup products on simplicial and cubical $n$–pseudomanifolds are presented. The author of [15] also defines L2 cubical Whitney forms and a cubical cup product that fits together with the wedge product of cubical Whitney forms.
In [16] cup products on cubical complexes have been applied to compute cohomology rings of 3D digital images (represented as cubical complexes $Q(I)$), thus simplify the combinatorial structure of $Q(I)$ and obtain a homeomorphic cellular complex $P(I)$ with fewer cells, which improves the computational efficiency. The authors of [16] think of the cup product as the way the $n$–dimensional holes obtained in homology are related to each other. It is known that two objects with non isomorphic cohomology rings are not homotopic.
Formulas for a diagonal approximation on a general polygon that are used to compute cup products on the cohomology of polygonal cellular complexes are presented in [17]. Their cup product uses a complex structure called the AT–model (see [17, Theorem 2]), and the coefficient of their cochains lie in the field $\mathbb{Z}_{2}$. Whereas the coefficients of our cochains lie in $\mathbb{R}$.
@@ 2 Results
We present a new cup product on polygonal 2–pseudomanifolds (Definition 15) and show that previously studied cup products on simplicial and cubical 2–pseudomanifolds are a special case of this more general polygonal cup product. We proceed by proving that the polygonal cup product satisfies the Leibniz product rule with discrete exterior derivative (Lemma 3) and it satisfies defining properties of an abstract cup product on cellular complexes (Theorem 4). Moreover, in Proposition 5 we show that the polygonal cup product shares several further properties with the wedge product of differential forms, such skew–commutativity. We thus use the expression polygonal cup product and polygonal wedge product interchangeably.
To properly define the polygonal cup product, we first review the appropriate discrete setting. In Section 3 we define orientable pseudomanifolds and the boundary operator. In Section 4 we formally introduce discrete differential forms as cochains and discrete exterior derivative as coboundary operator. Section 5 then contains our the actual results announced in the previous paragraph.
@@ 3 Discrete setting
The discrete domain over which our discrete wedge product operates is a general polygonal mesh. From the perspective of algebraic topology such a mesh corresponds to a 2–dimensional CW complex with some additional structures, concretely a 2–dimensional pseudomanifold. However, in this section we provide definitions without specifying the dimension.
A CW complex is a cellular complex, invented by J. H. C. Whitehead, and it is a more general complex that the simplicial one. According to [12, Paragraph 38], cellular complexes are more effective for computations of homology groups of given spaces. The following set of definitions were adapted from definitions in [12, Paragraph 38].
Definition 1.
@@
Next, we define pseudomanifold and introduce “orientability” for it through incidence numbers, which will also allow to define boundary homomorphism later on. For more details, see [14, Chapter IX].
Definition 3.
@@
If in the condition (2) of above definition, every $(n-1)$-dimensional cell is a face of exactly two $n$-cells, then we talk about pseudomanifold without boundary ([14, Chapter IX]). On the other hand, if there exist at least one $(n-1)$-cell which is incident with only one $n$-cell, we talk about pseudomanifold with boundary.
Definition 4.
@@ 4 Discrete Differential Forms and the Exterior Derivative
Similarly to the definition of chain complexes (Definition 8), we have the following:
Definition 11.
@@
A $q$–form $\alpha$ is said to be a closed form (cocycle) if $d\alpha=0$, and it is called an exact form (coboundary) if there exist a $(q-1)$–form $\beta$ such that $d\beta=\alpha$. Two closed $q$-forms are cohomologous if they differ by an exact $q$–form, i.e.,
@@
Note that each exact $q$-form $\alpha=d\beta$ is closed, since $d\alpha=dd\beta=0$.
@@ 5 Wedge and Cup Products
The wedge product exhibits the following properties, which we maintain also in the discrete setting except the associativity (compare with the later Proposition 5):
Proposition 1.
@@
For the proof see [18, Proposition 7.1.5 and Theorem 7.4.1].
@@
Whitney in [11, §9 of Appendix II] gives the following abstract definition of a cup product, which also appears in [15, Definition 2.3.3].
Definition 12.
@@
The interested reader is also referred to [14, Chapter XIII]. Whitney further asserts that thanks to the Leibniz rule being valid, we have the following:
Proposition 2.
@@
Comparing Proposition 1 with Definition 12 and Proposition 2, we affirm that both the wedge and the cup product, respectively, are bilinear operations that take as input two forms, resp. cochains, of arbitrary degree $p$ and $q$ and return a form, resp. cochain, of degree $p+q$. Moreover, they both satisfy the Leibniz rule. But a general cup products is guaranteed to be associative and skew–commutative only on cohomology. Fortunately, for our cup product on polygons we will recuperate the skew–commutativity for any discrete forms (not just closed ones). However, our polygonal cup product will be associative only on closed forms, see the Proposition 5.
Even though explicit formulas for computing cup product on a general cell complex are unknown, there are well–known explicit formulas for cup products on simplicial and cubical complexes [15]. We expand the set of known cup products with cup products on general polygons with coefficients in $\mathbb{R}$, which is one of our contributions.
@@ 5.1 Simplicial and cubical cup products
Arnold in [15, Definition 4.2.1] presents the definition of the Wilson’s cup product $\cup:C^{p}(K)\times C^{q}(K)\rightarrow C^{p+q}(K)$ given by
@@
where $W$ is the simplicial Whitney form (see [15, Definition 4.1.1]) and $c$ is an $(p+q)$–simplex.
Using the equation (4) we computed cup products on simplicial 2–dimensional complexes and state as the following definition:
Definition 13.
@@
$(\alpha^{1}\cup\beta^{1})(v_{0},v_{1},v_{2})$ $=$ $\frac{1}{6}\sum_{i=0}^{2}\alpha(v_{i},v_{i+1})(\beta(v_{i+1},v_{i+2})-\beta(v_{i-1},v_{i})),\;i\in\mathbb{Z}/3\mathbb{Z}.$
@@
The simplicial cup product satisfies the properties of Proposition 5, which we prove therein. Comparing the Proposition 5 with Propositions 2 and 1, we can see that our simplicial cup product is skew–commutative on any discrete forms, not only on cohomology as the generic cup product would require, but it is still not associative in general as the wedge product, only on cohomology.
In [15, Definition 3.2.1] the author defines a cubical cup product $\cup_{c}$, for which the product of two cochains agrees with the wedge product of their cubical Whitney forms (see [15, Section 3.2.2]), thus the Whitney map provides a connection of cubical cohomology with de Rham cohomology ([15, Theorem 3.2.12]). From [15, Definition 3.2.1] we computed the cup product on $k$–dimensional cubes for $0\leq k\leq 2$ and we present the formulas in the following definition:
Definition 14.
@@
$(\alpha^{1}\cup\beta^{1})(v_{0},v_{1},v_{2},v_{3})$ $=$ $\frac{1}{4}\sum_{i=0}^{3}\alpha(v_{i},v_{i+1})(\beta(v_{i+1},v_{i+2})-\beta(v_{i-1},v_{i})),\;i\in\mathbb{Z}/4\mathbb{Z}.$
@@ 5.2 The cup product on polygons
As advertised earlier, we can extend the definitions of the discrete wedge product to any $p$–polygon such that it still satisfies the defining properties of a cup product (Definition 12) and therefore automatically also the properties of Proposition 2. Moreover, we show that our polygonal cup product is a skew–commutative bilinear operation on any polygonal forms, not just on closed ones, viz. the Proposition 5.
@@
It is easy to see that the simplicial and cubical cup products (Definitions 13 and 14) are just special cases of the product defined above.
In Theorem 4 we prove that the polygonal cup product actually is a cup product, that is, we prove that it satisfies the defining properties of a cup product (Definition 12). But we first prove, in the following lemma, that it obeys the Leibniz rule with the discrete derivative.
Lemma 3.
The discrete wedge product from Definition 15 satisfies the Leibniz rule with the discrete derivative, that is:
@@
$(d_{0}\alpha\cup\beta+\alpha\cup d_{0}\beta)(e)$ $=(\alpha(v_{j})-\alpha(v_{i}))\frac{\beta(v_{i})+\beta(v_{j})}{2}+\frac{\alpha(v_{i})+\alpha(v_{j})}{2}(\beta(v_{j})-\beta(v_{i}))$
@@
$=\alpha(v_{j})\beta(v_{j})-\alpha(v_{i})\beta(v_{i})=(\alpha\cup\beta)(v_{j})-(\alpha\cup\beta)(v_{i})$
@@
$=$ $\;n\beta_{k}-2(\beta_{k}+\beta_{k+1}+\beta_{k+2}+\beta_{k+3}+\dots+\beta_{k+\frac{n}{2}-3}+\beta_{k+\frac{n}{2}-2}+\beta_{k+\frac{n}{2}-1}),$
@@
$=$ $\;n\beta_{k-1}-2(\beta_{k-1}+\beta_{k-2}+\beta_{k-3}+\dots+\beta_{k+\frac{n}{2}+2}+\beta_{k+\frac{n}{2}+1}+\beta_{k+\frac{n}{2}}).$
@@
$=$ $\;n\beta_{k}-2(\beta_{k}+\beta_{k+1}+\beta_{k+2}+\beta_{k+3}+\dots+\beta_{k+\frac{n-1}{2}-2}+\beta_{k+\frac{n-1}{2}-1})-\beta_{k+\frac{n-1}{2}},$
@@
$=$ $\;n\beta_{k-1}-2(\beta_{k+\frac{n-1}{2}+1}+\beta_{k+\frac{n-1}{2}+2}+\dots+\beta_{k-3}+\beta_{k-2}+\beta_{k-1})-\beta_{k+\frac{n-1}{2}}.$
@@
Theorem 4.
The polygonal cup product from Definition 15 is a cup product, i.e., it satisfies the Definition 12.
@@
It also satisfies item 1 of Definition 12 – the result of product of two forms is located on a face that is incident to both the elements where the two original discrete forms reside.
Due to Lemma 3, our product satisfies item 2, the Leibniz rule.
We will now prove that item 3 of Definition 12 also holds. A pseudomanifold $K$ as we defined it is a connected cell complex, therefore we have to show that there exist a real number $\xi$ such that $I^{0}\cup\alpha^{p}=\alpha^{p}\cup I^{0}=\xi\alpha^{p}$, where $I^{0}$ is the constant 0–form that takes value 1 on the vertices of $K$. Using the equations of Definition 15, we can thus write:
@@
In the following proposition, which can be thought of as an analog of Proposition 1, we summarize and prove further properties of our polygonal cup product.
Proposition 5.
Let $K$ be a 2–pseudomanifold, $\alpha^{k}\in C^{k}(K)$, $\beta^{l}\in C^{l}(K)$, and $\gamma^{m}\in C^{m}(K)$. The polygonal cup product from Definition 15 satisfies these properties:
@@
From Theorem 4 we know our product is a bilinear operation satisfying the Leibniz rule with discrete exterior derivative. We now show that it is also skew–commutative on any discrete forms and associative in the declared fashion.
With respect to the announced form of associativity, in the case of the product of three 0–forms we get to multiplying three real numbers on each vertex, which is indeed an associative operation. For the other degrees, we have learned in the proof of Theorem 4 that if $K$ is connected and a 0–form $\alpha^{0}$ is constant (which is equivalent to closed), that is $\alpha^{0}(v_{i})=\rho\in\mathbb{R}\;\forall v_{i}\in K$, then the cup product of $\alpha^{0}$ with any $q$–form $\omega^{q}$ is equal to multiplying $\omega^{q}$ by the real number $\rho$. This fact yields:
@@
$+\alpha(-a)\beta(0)+\alpha(1-a)\beta(1)+\alpha(2-a)\beta(2)+\cdots+\alpha(n-1)\beta(n-1+a)$
@@
$+\alpha(a)\beta(0)+\alpha(1+a)\beta(1)+\alpha(2+a)\beta(2)+\cdots+\alpha(n-1)\beta(n-1-a)$
@@
$(\alpha^{1}\cup\beta^{1})(f)=-\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}\bigg{(}\frac{1}{2}-\frac{a}{n}\bigg{)}\sum_{i=0}^{n-1}\beta(i)(\alpha((i+a)\%n)-\alpha((i-a)\%n))=-(\beta^{1}\cup\alpha^{1})(f).$

### v3

@@ Abstract
Discrete exterior calculus offers a coordinate–free discretization of exterior calculus especially suited for computations on curved spaces. In this work, we present a wedge product on 2–dimensional pseudomanifolds, whose faces are any polygons. We prove that this polygonal wedge product is compatible with the discrete exterior derivative in the sense that it satisfies the Leibniz product rule. We thus extend previously studied discretizations of wedge products from simplicial or quadrilateral meshes to surface meshes whose faces are arbitrary simple polygons. We also prove that our discrete wedge product corresponds to a cup product of cochains on 2–pseudomanifolds. By rigorously justifying our construction we add another piece to ever evolving discrete versions of exterior calculus.
@@
Discrete wedge product , Discrete exterior calculus , Polygonal wedge product , Polygonal cup product , Discrete differential forms , Polygonal exterior calculus.
MSC:
65N22 , 57Z20 , 57-08 , 53Z30
[…]
organization=Dept. of Applied Mathematics, Charles University, addressline=Ke Karlovu 3, city=Praha 2, postcode=121 12, country=Czech Republic
@@ 1 Introduction
Discrete versions of exterior calculus on surface meshes have gained increased attention in the last few decades, see Hirani (2003); Arnold (2018); Desbrun et al. (2008); Ptackova and Velho (2021) and references therein. Discete exterior calculus (DEC) and finite element exterior calculus (FEEC) have found applications in computational fluid dynamics (Mohamed et al., 2016; Nitschke et al., 2017), computational electromagnetism (Zhang et al., 2021), and computer graphics (Desbrun et al., 2003; Ptackova and Velho, 2021). See also references in Schubel et al. (2024) for more applications.
In this paper, we prove that the wedge product on general polygonal meshes, first introduced in Ptackova and Velho (2017) and later used in applications in Ptackova and Velho (2021), satisfies the Leibniz product rule with discrete exterior derivative (coboundary operator). Furthermore, we show that it corresponds to a cup product of cochains on orientable 2–dimensional pseudomanifolds. We thus partially extend the findings of Schubel et al. (2024) from simplicial to general polygonal setting.
On smooth manifolds, the wedge product allows for building higher degree forms from lower degree ones. Similarly a cup product is a product of two cochains of arbitrary degree $p$ and $q$ that returns a cochain of degree $p+q$ located on $(p+q)$–dimensional cells. The cup product was introduced by J. W. Alexander, E. Čech, and H. Whitney (Whitney, 1957) in 1930’s and it became a well–studied notion in algebraic topology, mainly in the simplicial setting (Munkres, 1984; Fenn, 1983). Later, the cup product was extended also to $n$–cubes (Massey, 1991). In (Arnold, 2012) various cup products on simplicial and cubical $n$–pseudomanifolds are presented. Arnold (2012) also defines L2 cubical Whitney forms and a cubical cup product that fit together with the wedge product of cubical Whitney forms.
In (Gonzalez-Diaz et al., 2011a) cup products on cubical complexes have been applied to compute cohomology rings of 3D digital images (represented as cubical complexes $Q(I)$), thus simplify the combinatorial structure of $Q(I)$ and obtain a homeomorphic cellular complex $P(I)$ with fewer cells, which improves the computational efficiency. Gonzalez-Diaz et al. (2011a) think of the cup product as the way the $n$–dimensional holes obtained in homology are related to each other. It is known that two objects with non isomorphic cohomology rings are not homotopic.
Formulas for a diagonal approximation on a general polygon that are used to compute cup products on the cohomology of polygonal cellular complexes are presented in (Gonzalez-Diaz et al., 2011b). Their cup product uses a complex structure called the AT–model (Gonzalez-Diaz et al., 2011b, Theorem 2), and the coefficient of their cochains lie in the field $\mathbb{Z}_{2}$. Whereas the coefficients of our cochains lie in $\mathbb{R}$.
@@ 2 Results
We present a new cup product on polygonal 2–pseudomanifolds (Definition 5.4) and show that previously studied cup products on simplicial and cubical 2–pseudomanifolds are a special case of this more general polygonal cup product. We proceed by proving that the polygonal cup product satisfies the Leibniz product rule with discrete exterior derivative (Lemma 5.1) and it satisfies defining properties of an abstract cup product on cellular complexes (Theorem 5.1). Moreover, in Proposition 5.3 we show that the polygonal cup product shares several further properties with the wedge product of differential forms, such skew–commutativity. We thus use the terms polygonal cup product and polygonal wedge product interchangeably.
To properly define the polygonal cup product, we first review the appropriate discrete setting in Section 3, concretely we define orientable pseudomanifolds, which is the domain of our interest, and we revise the boundary operator of chains of cells. In Section 4 we formally introduce discrete differential forms as cochains and discrete exterior derivative as coboundary operator. Section 5 then contains our the actual results announced in the previous paragraph.
@@ 3 Discrete setting
The discrete domain over which our discrete wedge product operates is a general polygonal mesh, i.e., a surface mesh whose faces are arbitrary simple polygons. From the perspective of algebraic topology such a mesh corresponds to a 2–dimensional CW complex with some additional structures: a 2–dimensional pseudomanifold. In this section we briefly revise several definitions that ultimately leads to the definition of an orientable pseudomanifold and the boundary operator.
A CW complex is a cellular complex, invented by J. H. C. Whitehead, and it is a more general complex that the simplicial one. According to (Munkres, 1984, Paragraph 38), cellular complexes are more effective for computations of homology groups of given spaces. The following set of definitions were adapted from definitions in (Munkres, 1984, Paragraph 38).
Definition 3.1.
@@
Next, we define pseudomanifold and introduce its “orientability” through incidence numbers, which will also allow to define boundary homomorphism later on. For more details, see (Massey, 1991, Chapter IX).
Definition 3.3.
@@
If in the condition (2) of above definition, every $(n-1)$-dimensional cell is a face of exactly two $n$-cells, then we talk about pseudomanifold without boundary ((Massey, 1991, Chapter IX)). On the other hand, if there exist at least one $(n-1)$-cell which is incident with only one $n$-cell, we talk about pseudomanifold with boundary.
Definition 3.4.
@@ 4 Discrete Differential Forms and the Exterior Derivative
Similarly to the definition of chain complexes (Definition 3.8), we have the following:
Definition 4.3.
@@
A $q$–form $\alpha$ is said to be a closed form (cocycle) if $d\alpha=0$, and it is called an exact form (coboundary) if there exist a $(q-1)$–form $\beta$ such that $d\beta=\alpha$.
Note that each exact $q$-form $\alpha=d\beta$ is closed, since $d\alpha=dd\beta=0$. Two closed $q$-forms are cohomologous if they differ by an exact $q$–form. We denote cohomology classes by $H^{\ast}$:
@@ 5 Wedge and Cup Products
The wedge product exhibits the following properties, which we maintain also in the discrete setting except the associativity (compare with the later Proposition 5.3):
Proposition 5.1.
@@
For the proof see (Abraham et al., 1988, Proposition 7.1.5 and Theorem 7.4.1).
@@
Whitney in (Whitney, 1957, §9 of Appendix II) gives the following abstract definition of a cup product, which also appears in (Arnold, 2012, Definition 2.3.3).
Definition 5.1.
@@
The interested reader is also referred to (Massey, 1991, Chapter XIII). Whitney further asserts that thanks to the Leibniz rule being valid, we have the following:
Proposition 5.2.
@@
Comparing Proposition 5.1 with Definition 5.1 and Proposition 5.2, we affirm that both the wedge and the cup product, respectively, are bilinear operations that take as input two forms, resp. cochains, of arbitrary degree $p$ and $q$ and return a form, resp. cochain, of degree $p+q$. Moreover, they both satisfy the Leibniz rule. But a general cup products is guaranteed to be associative and skew–commutative only on cohomology. Fortunately, for our cup product on polygons we will recuperate the skew–commutativity for any discrete forms (not just closed ones). However, our polygonal cup product will be associative only on closed forms, see the Proposition 5.3.
Even though explicit formulas for computing cup product on a general cell complex are unknown, there are well–known explicit formulas for cup products on simplicial and cubical complexes (Arnold, 2012). We expand the set of known cup products with cup products on general polygons with coefficients in $\mathbb{R}$, which is one of our contributions.
@@ 5.1 Simplicial and cubical cup products
Let us first recall the definition of simplicial Whitney forms. Let $c=(v_{0},\dots,v_{p})$ be a $p$–simplex with barycentric coordinates $x_{0},\dots,x_{p}$. Thus $0\leq x_{j}\leq 1\;\forall j$ and $\sum_{i=0}^{p}=1$. The simplicial Whitney $p$–form corresponding to the simplex $c$ is
$W(c)=p!\sum_{i=0}^{p}(-1)^{i}x_{i}dx_{0}\wedge\dots\wedge\widehat{dx_{i}}\wedge\dots\wedge dx_{p},$
where $\;{}\widehat{}\;$ indicates a term omitted from the product. This definition extends linearly to define Whitney forms on all of $C^{\ast}(K)$.
Arnold in (Arnold, 2012, Definition 4.2.1) presents the following definition of a cup product on simplices, which she calls the Wilson’s cup product, $\cup:C^{p}(K)\times C^{q}(K)\rightarrow C^{p+q}(K)$ given by
@@
where $c$ is an $(p+q)$–simplex and $W$ is the simplicial Whitney form.
Even though the above cup product has appeared already in (Whitney, 1957), it has been studied by Scott Wilson in (Wilson, 2007), where he proved that the cup product converges to wedge product of differential forms (Wilson, 2007, Theorem 5.4). Furthermore, Arnold claims that Wilson’s work was the motivation for the majority of results presented in (Arnold, 2012), including the definition of cubical cup product presented in the next paragraphs.
Using the equation (4) we compute cup products on 2–dimensional simplicial complexes and state them as the following definition:
Definition 5.2.
@@
$(\alpha^{1}\cup\beta^{1})(v_{0},v_{1},v_{2})$ $=$ $\frac{1}{6}\sum_{i=0}^{2}\alpha(v_{i},v_{i+1})(\beta(v_{i+1},v_{i+2})-\beta(v_{i-1},v_{i})),$
where the indices are taken to be modulo 3.
@@
The simplicial cup product satisfies the properties of Proposition 5.3, which we prove therein. Comparing the Proposition 5.3 with Propositions 5.2 and 5.1, we can see that our simplicial cup product is skew–commutative on any discrete forms, not only on cohomology as the generic cup product would require, but it is still not associative in general as the wedge product, only on cohomology.
In (Arnold, 2012, Definition 3.2.1) the author defines a cubical cup product $\cup_{c}$, for which the product of two cochains agrees with the wedge product of their cubical Whitney forms (see (Arnold, 2012, Section 3.2.2)), thus the Whitney map provides a connection of cubical cohomology with de Rham cohomology ((Arnold, 2012, Theorem 3.2.12)). From (Arnold, 2012, Definition 3.2.1) we computed the cup product on $k$–dimensional cubes for $0\leq k\leq 2$ and we present the formulas in the following definition:
Definition 5.3.
@@
$(\alpha^{1}\cup\beta^{1})(v_{0},v_{1},v_{2},v_{3})$ $=$ $\frac{1}{4}\sum_{i=0}^{3}\alpha(v_{i},v_{i+1})(\beta(v_{i+1},v_{i+2})-\beta(v_{i-1},v_{i})),$
where the indices are taken to be modulo 4.
@@ 5.2 The cup product on polygons
As advertised earlier, we can extend the definitions of the discrete wedge product to any $p$–polygon such that it still satisfies the defining properties of a cup product (Definition 5.1) and therefore automatically also the properties of Proposition 5.2. Moreover, we show that our polygonal cup product is a skew–commutative bilinear operation on any polygonal forms, not just on closed ones, viz. the Proposition 5.3.
@@
It is easy to see that the simplicial and cubical cup products (Definitions 5.2 and 5.3) are just special cases of the product defined above.
In Theorem 5.1 we prove that the polygonal cup product actually is a cup product, that is, we prove that it satisfies the defining properties of a cup product (Definition 5.1). But we first prove, in the following lemma, that it obeys the Leibniz rule with the discrete derivative.
Lemma 5.1.
The discrete wedge product from Definition 5.4 satisfies the Leibniz rule with the discrete derivative, that is:
@@
$(d_{0}\alpha\cup\beta+\alpha\cup d_{0}\beta)(e)$ $=(\alpha(v_{j})-\alpha(v_{i}))\frac{\beta(v_{i})+\beta(v_{j})}{2}$
$+\frac{\alpha(v_{i})+\alpha(v_{j})}{2}(\beta(v_{j})-\beta(v_{i}))$
@@
$=\alpha(v_{j})\beta(v_{j})-\alpha(v_{i})\beta(v_{i})$
$=(\alpha\cup\beta)(v_{j})-(\alpha\cup\beta)(v_{i})$
@@
$=$ $\;n\beta_{k}-2(\beta_{k}+\beta_{k+1}+\beta_{k+2}+\beta_{k+3}+\dots$
$+\beta_{k+\frac{n}{2}-3}+\beta_{k+\frac{n}{2}-2}+\beta_{k+\frac{n}{2}-1}),$
@@
$=$ $\;n\beta_{k-1}-2(\beta_{k-1}+\beta_{k-2}+\beta_{k-3}+\dots$
$+\beta_{k+\frac{n}{2}+2}+\beta_{k+\frac{n}{2}+1}+\beta_{k+\frac{n}{2}}).$
@@
$=$ $\;n\beta_{k}-2(\beta_{k}+\beta_{k+1}+\beta_{k+2}+\beta_{k+3}+\dots+\beta_{k+\frac{n-1}{2}-2}+\beta_{k+\frac{n-1}{2}-1})$
$-\beta_{k+\frac{n-1}{2}},$
@@
$=$ $\;n\beta_{k-1}-2(\beta_{k+\frac{n-1}{2}+1}+\beta_{k+\frac{n-1}{2}+2}+\dots+\beta_{k-3}+\beta_{k-2}+\beta_{k-1})$
$-\beta_{k+\frac{n-1}{2}}.$
@@
Theorem 5.1.
The polygonal cup product from Definition 5.4 is a cup product, i.e., it satisfies the Definition 5.1.
@@
It also satisfies item 1 of Definition 5.1 – the result of product of two forms is located on a face that is incident to both the elements where the two original discrete forms reside.
Due to Lemma 5.1, our product satisfies item 2, the Leibniz rule.
We will now prove that item 3 of Definition 5.1 also holds. A pseudomanifold $K$ as we defined it is a connected cell complex, therefore we have to show that there exist a real number $\xi$ such that $I^{0}\cup\alpha^{p}=\alpha^{p}\cup I^{0}=\xi\alpha^{p}$, where $I^{0}$ is the constant 0–form that takes value 1 on the vertices of $K$. Using the equations of Definition 5.4, we can thus write:
@@
In the following proposition, which can be thought of as an analog of Proposition 5.1, we summarize and prove further properties of our polygonal cup product.
Proposition 5.3.
Let $K$ be a 2–pseudomanifold, $\alpha^{k}\in C^{k}(K)$, $\beta^{l}\in C^{l}(K)$, and $\gamma^{m}\in C^{m}(K)$. The polygonal cup product from Definition 5.4 satisfies these properties:
@@
From Theorem 5.1 we know our product is a bilinear operation satisfying the Leibniz rule with discrete exterior derivative. We now show that it is also skew–commutative on any discrete forms and associative in the declared fashion.
With respect to the announced form of associativity, in the case of the product of three 0–forms we get to multiplying three real numbers on each vertex, which is indeed an associative operation. For the other degrees, we have learned in the proof of Theorem 5.1 that if $K$ is connected and a 0–form $\alpha^{0}$ is constant (which is equivalent to closed), that is $\alpha^{0}(v_{i})=\rho\in\mathbb{R}\;\forall v_{i}\in K$, then the cup product of $\alpha^{0}$ with any $q$–form $\omega^{q}$ is equal to multiplying $\omega^{q}$ by the real number $\rho$. This fact yields:
@@
$+\alpha(-a)\beta(0)+\alpha(1-a)\beta(1)+\alpha(2-a)\beta(2)+\cdots$
$+\alpha(n-1)\beta(n-1+a)$
@@
$+\alpha(a)\beta(0)+\alpha(1+a)\beta(1)+\alpha(2+a)\beta(2)+\cdots$
$+\alpha(n-1)\beta(n-1-a)$
@@
$(\alpha^{1}\cup\beta^{1})(f)$ $=-\sum_{a=1}^{\lfloor\frac{n-1}{2}\rfloor}\bigg{(}\frac{1}{2}-\frac{a}{n}\bigg{)}\sum_{i=0}^{n-1}\beta(i)\big{(}\alpha((i+a)\%n)-\alpha((i-a)\%n)\big{)}$
$=-(\beta^{1}\cup\alpha^{1})(f).$